TheRiddlersRiddle:

What are the dimensions if the poster is enlarged by a factor of five over two ? A rectangle is shown. The length of the rectangle is labeled 12 feet. The width of the rectangle is labeled 3 feet.

Questionnare:

Okay, I am in ur class so uhh lmao If you enlarge something by factoring, you multiply the measurements. 12 times 5/2 = 30 (5/2 = 2.5) 3 times 5/2 = 7.5 The new length of the rectangle is 30 feet squared. The width of the rectangle is now 7.5 feet squared.